OpenStudy (anonymous):

Which factor can you cancel from the numerator and denominator to simplify 8-2x/3x-12??

OpenStudy (anonymous):

(8 - 2x)/(3x -12) = 2(4 - x)/(3(x - 4)) = -2(x -4)/(3(x - 4)) = -2/3 (4 - x) = -(x - 4) so it requires some manipulation to get the common factor clearly visible